OpenStudy (anonymous):

express 120° 15' 36" in decimal degree form

OpenStudy (anonymous):

Use the relationships 1º = 60’ and 1’ = 60”. 120º 15’ 36” Convert the fractions of a degree to decimals using dimensional analysis, then add to 120º. 15’ * (1º/60’) = 15/60 = 0.25º, and 36” * (1’/60”) * (1º/60’) = 36/3600 = 0.01º Now add: 120º 15’ 36” = 120º + 0.25º + 0.01º = 120.26º